# Yuyang Tao, Shufei Ge

### v1

@@ Abstract
Motivation: The Mapper algorithm is an essential tool to explore shape of data in topology data analysis. With a dataset as an input, the Mapper algorithm outputs a graph representing the topological features of the whole dataset. This graph is often regarded as an approximation of a reeb graph of data. The classic Mapper algorithm uses fixed interval lengths and overlapping ratios, which might fail to reveal subtle features of data, especially when the underlying structure is complex.
Results: In this work, we introduce a distribution guided Mapper algorithm named D-Mapper, that utilizes the property of the probability model and data intrinsic characteristics to generate density guided covers and provides enhanced topological features. Our proposed algorithm is a probabilistic model-based approach, which could serve as an alternative to non-prababilistic ones. Moreover, we introduce a metric accounting for both the quality of overlap clustering and extended persistence homology to measure the performance of Mapper type algorithm. Our numerical experiments indicate that the D-Mapper outperforms the classical Mapper algorithm in various scenarios. We also apply the D-Mapper to a SARS-COV-2 coronavirus RNA sequences dataset to explore the topological structure of different virus variants. The results indicate that the D-Mapper algorithm can reveal both vertical and horizontal evolution processes of the viruses.
Availability: Our package is available at https://github.com/ShufeiGe/D-Mapper.
Contact: geshf@shanghaitech.edu.cn
@@ 1 Introduction
Topology data analysis (TDA) is a specific area of investigating data shapes based on topological theory. Topology is a powerful tool to study the shape of objects by finding invariants that remain unchanged under continuous deformations. These invariants can reflect objects’ intrinsic features. TDA attempts to identify and utilize these topology invariants in data analysis. The Mapper algorithm is a simple but powerful TDA tool for visualizing and clustering data. The algorithm constructs a graph in which each vertex represents a cluster, and each edge represents the two adjacent clusters that share some elements. The colour and size of vertices are defined by users and can provide more information. The size of vertices and the colour depth can represent the average value and the number of elements in the cluster. The Mapper algorithm was first proposed in singhTopologicalMethodsAnalysis and has been applied in various domains, such as social media almgrenMiningSocialMedia2017 , fraud detection mitraExperimentsFraudDetection2021 , and evolutionary computation todaVisualizationClusteringGraph2022 . The Mapper algorithm is especially suitable for biology data, which are usually complex and high-dimensional skafTopologicalDataAnalysis2022 . For example, the Mapper algorithm was applied to breast cancer transcriptional microarray data and successfully identified two subgroups with $100\%$ survival rate nicolauTopologyBasedData2011 . It is also used to analyze transcriptional programs that control cellular lineage commitment and differentiation during development, and the proposed scTDA identified four transient states over time rizviSinglecellTopologicalRNAseq2017 . Exploring the conformation space of proteins is another important task in computational biology, and the Mapper algorithm has been successfully applied to analyze intermediate conformations, the results are closely consistent with experimental findings dehghanpoorUsingTopologicalData .
@@
The classic Mapper algorithm requires to select a filter function $f:X\rightarrow\mathbb{R}^{n}$ that projects data $X$ onto the Euclidean space $\mathbb{R}^{n}$ and a set of open cover $\mathcal{U}$ on the projection data. Different filter functions and open covers may result in different outputs, so it is necessary to select them carefully. Improper filter functions or open covers may not accurately reveal data shapes, resulting in poor overlap clustering. A filter function is chosen primarily based on features of interest and also depends on specific applications. Determining the optimal parameters of the model, such as the overlapping rates and interval lengths, usually requires extensive manual tuning and experiments. As pointed out in carriereStatisticalAnalysisParameter2017 , cover choices lead to a broad range of possible Mapper outputs, resulting in many different graph depictions and clustering results. Unreasonable choices may cause the disappearance of certain topological features.
@@
Many works were proposed to improve the performance of the classic Mapper. For example, dlotkoBallMapperShape2019 proposed to generate covers on the dataset by constructing a set of balls directly. This method saved trouble of choosing filter function. buiFMapperFuzzyMapper2020 used fuzzy clustering algorithm to generate covers to choose cover automatically with random overlap ratios. To achieve adaptive cover construction, a information criteria was developed based on X-means algorithm to generate adaptive covers chalapathiAdaptiveCoversMapper2021 .
@@
In this work, alternatively, we propose a distribution-guided Mapper algorithm to relax the restriction of regular covers and overlapping. Our proposed algorithm utilizes the property of the probability model and data intrinsic characteristics to generate distribution guided covers and provides enhanced topological features. Our proposed algorithm is a probabilistic model-based approach, which could serve as an alternative to non-probabilistic ones. Unlike the classic Mapper algorithm, our proposed D-Mapper does not rely on predefined static overlap rates and interval lengths, but instead we fit projected data to a mixture distribution model and generate flexible covers automatically. Moreover, the performance of Mapper algorithm is hard to evaluate due to the complexity of output graph structure, most of literature only evaluate from the clustering perspective buiFMapperFuzzyMapper2020 ; chalapathiAdaptiveCoversMapper2021 , however these method overlooks the quality of topological structure. In this paper, We introduce a metric that can quantitatively and objectively evaluate the performance of Mapper algorithm from both clustering and topological aspects. We use the extended persistence homology as a tool to capture topological signatures, then we distinguish noise and signal on the diagram by constructing confidence set on the diagram, the signal rate can be a metric to evaluate the quality of topological structure. All the simulation and application studies indicate that the D-Mapper algorithm outperforms the classic Mapper algorithm. The reminder of the paper is organized as follows. We briefly review the classic Mapper algorithm and introduce our D-Mapper and evaluation metric in the Materials and methods. Simulation section focuses on the comparison of the classic Mapper and the D-Mapper on simulated datasets. In Application section, we applied D-Mapper algorithm on a real world SARS-COV-2 RNA sequences dataset. We summarize our work and discuss the future work in the conclusion section.
2 Materials and methods
@@ 2.1 Basic notions
Before we introduce our method, we first briefly review the essential background of the Mapper algorithm. The theoretical foundation of the Mapper algorithm is the Nerve theorem which guarantees the nerves produced by a cover on a topology space $X$ is homomorphic equivalent to that space. We introduce the following notions to describe Nerve theorem carlssonTopologyData2009 ; chazalIntroductionTopologicalData2021 .
@@
Given a set $P=\{p_{0},...,p_{k}\}\subset\mathbb{R}^{n}$ of k+1 affinely independent points, a $k$-dimensional simplex $\sigma$, or $k$-simplex for short, spanned by $P$ is the set of convex combinations such that:
@@
b) The intersection of any two simplices of $K$ is either empty or the common face of the two.
The union of the simplices of $K$ constitutes the underlying space of $K$, denoted as $|K|$, which inherits from the topology of $\mathbb{R}^{n}$. Thus the geometric simplicial complex can also be regarded as a topology space.
@@
Suppose $\mathcal{U}=(u_{i}),i\in I$ is a collection of open subset of a topological space $X$, then we say $\mathcal{U}$ is an open cover of $X$ if $X\subset\bigcup_{i\in I}u_{i}$.
Given an open cover of topological space $X$, $\mathcal{U}=(u_{i})$, the nerve of $\mathcal{U}$ is an abstract simplicial complex C$(\mathcal{U})$ with vertex set $\mathcal{U}$. With these definitions, we can introduce the most important theorem in constructing Mapper.
@@
Let $\mathcal{U}=(u_{i}),i\in I$ be a cover of a topological space $X$ by open sets such that the intersection of any sub-collection of the $u_{i}$’s is either empty or contractible. Then, $X$ and the nerve C($\mathcal{U}$) are homotopy equivalent.
This theorem allows us to map the topology of continuous into abstract combinatorial structures by building a nerve complex. It bridges the gap between continuous space and its discrete representation. Many TDA methods, including the Mapper algorithm, are built based on this crucial theorem.
@@ 2.2 Mapper algorithm
Mapper algorithm is an important tool to construct the discrete version of reeb graphs which encode connected information of the support manifold deyTopologicalAnalysisNerves2017 . Algorithm 1 depicts the classic Mapper algorithm. Firstly, the original data is projected onto a real line by a user-specified filter function (Algorithm 1, line 1). To construct a cover $\mathcal{U}$ on the projected data, the number of components $n$ and overlap ratio $p$ should be chosen carefully (Algorithm 1, line 2). The improper choice of these parameters may lead to failure estimation of the shape of data. With these two parameters, a cover with $n$ regular intervals can be obtained (Algorithm 1, line 3). Then, the inverse image of the cover $\mathcal{U}$ is achieved on the original data, generating some hypercubes. This process is often called pulling back (Algorithm 1, line 4). Finally, clustering data points within each hypercubes. Each cluster corresponds to a vertex in the Mapper graph. If two vertexes share any elements, an edge is added between these two vertices (Algorithm 1, line 5).
@@
1:Choose a proper filter function $f$ to project data on the real line, $f:X\rightarrow\mathbb{R}$.
2:Choose a component number $n$ and overlap percentage ratio $p$.
@@
4:Pull back the intervals of projected data, $f^{-1}(\mathcal{U})$.
@@
The Mapper algorithm is powerful for exploring and visualizing data, while the selection of parameters $n,~{}p$ involves extensive manual tuning and the algorithm is sensitive to these parameters. In addition, the Mapper algorithm’s flexibility is often restricted by regular intervals and fixed overlap ratios, which may hinder the discovery of complex data structures.
@@ 2.3 D-Mapper
The restriction of regular intervals with fixed overlap ratios is one of the major limitations of the Mapper algorithm. We propose a distribution-guided Mapper algorithm to generate flexible covers to reflect the underlying data structures. Our proposed method automatically chooses the overlap ratios based on the distribution of the projection data and produces more flexible covers to reveal the data shapes more accurately. Our experiments indicate that our method produces better results than the classic Mapper algorithm in terms of clustering. The key idea of our algorithm is to fit projected data with a mixture model. Each component in the mixture model can be viewed as a cluster, and the probability (likelihood) of each data point assigned to each cluster can be explicitly calculated. Once we get the mixture distribution of projected data, we can create intervals base on the distribution in many ways.
Here we introduce a simply way to construct intervals based on the quantile $\alpha$ of the probability distribution. With proper selection, the $\alpha$ quantile interval of each component of the mixture model can automatically produce some overlaps, Figure 1 shows the idea of naturally producing intervals by quantile $\alpha$. This attribute provides a natural scheme for constructing flexible covers on projected data. The specific procedures are as follows:
@@
Figure 1: An illustration of intervals produced by D-Mapper. The deep blue line is the probability density function of each component in the GMM. The shallow blue dashed line is the probability density function of the GMM. The orange lines are intervals that produced naturally by symmetry quantile $\alpha$. (a) The quantile $\alpha=0.01$, each interval has pairwise overlap. (b) The quantile $\alpha=0.1$, there is a gap between the first and second interval. The quantile $\alpha$ can control the overlap of intervals, therefore this parameter should be chosen carefully.
In both our simulation and real data experiments, we implement the Gaussian mixture models (GMM) to fit the data due to its simplicity and flexibility. The model inference is done via the expectation maximization (EM) algorithm 2009The . Other mixture models or any distribution with multiple modes can be used as alternatives to the GMM.
By incorporating a mixture model into the Mapper algorithm, it can produce flexible intervals and overlaps. We call this distribution-guided Mapper algorithm as D-Mapper algorithm. With these flexible intervals, we can construct the output graph similar to the classic Mapper algorithm: pulling back these intervals onto the original data space, then applying clustering algorithm within each intervals separately, all sub-groups constitute the vertices of the nerve, and adding an edge between two vertices if there are any shared data points. Algorithm 2 gives details of the D-Mapper.
@@
1:Choose a proper filter function $f$ to project data on the real line,$f:X\rightarrow\mathbb{R}$.
@@
3:Fit projected data to a mixture model.
@@
10:Pull back the intervals of projected data, $f^{-1}(\mathcal{U})$.
11:Clustering on the refined cover and build the nerve by clustering result.
The quantile $\alpha$ controls the overlap ratios: larger $\alpha$ leads to lower overlap ratios. Distributions of components in mixture models are often heterogeneous, result in irregular quantile intervals. Proper $\alpha$ selection is crucial to ensure all points are covered.
The Mapper algorithms usually require pairwise overlap (i.e. each interval has overlap with its neighbours)deyTopologicalAnalysisNerves2017 ; dlotkoBallMapperShape2019 ; buiFMapperFuzzyMapper2020 ; chalapathiAdaptiveCoversMapper2021 . For D-Mapper, we also want to preserve the pairwise overlap property (Figure 1 (a)). However, larger $\alpha$ may result in disjoint intervals, Figure 1 (b) gives an example of disjoint intervals caused by improper large $\alpha$. Thus, in this work, we propose a method to find the upper bound of $\alpha$ that guarantees the pairwise overlap property. With this upper bound, we also greatly reduce the range of $\alpha$ to make parameter selection easier. Denote $F_{i}^{-1}$ the inverse of the cumulative density function of the $i$th ordered component, thus the $i$th interval $u_{i}$ is given by $[F_{i}^{-1}(\alpha/2),F_{i}^{-1}(1-\alpha/2)]$, $i=1,\ldots,n$. The idea of this method is to find the upper bound $\alpha^{\prime}$, that guarantee the intersection of each paired neighboring intervals is not empty, $u_{i}\cap u_{i+1}\neq\emptyset$, i.e., $F_{i}^{-1}(1-\alpha/2)\geq F_{i+1}^{-1}(\alpha/2)$, $i=1,2,\ldots,n-1$. Algorithm 3 describes how to find the upper bound of $\alpha$.
@@
6:     if $*\alpha_{i}^{\prime}\geq\alpha^{a}*$ then
@@
Notice that although pairwise overlap is a good property for a Mapper graph, this property is not necessary for all situations. One can construct a good Mapper graph with some paired neighbouring intervals being non-overlap as long as covers can cover all data points. We add one threshold parameter $\alpha^{*}$ to allow for disjoint paired neighbouring intervals (lines 6-10 of Algorithm 3). In practice, we suggest to set $\alpha^{*}=0.005$. As shown in the simulation section, the example of the two disjoint circles does not overlap in pairs (Figure 5 (a) ), since there are no points between the two circles, disjoint paired neighbouring intervals should be presented in the corresponding Mapper output graph.

### v2

# A distribution-guided Mapper algorithm

@@ Abstract
Background: The Mapper algorithm is an essential tool to explore the shape of data in topology data analysis. With a dataset as an input, the Mapper algorithm outputs a graph representing the topological features of the whole dataset. This graph is often regarded as an approximation of a Reeb graph of data. The classic Mapper algorithm uses fixed interval lengths and overlapping ratios, which might fail to reveal subtle features of data, especially when the underlying structure is complex.
Results: In this work, we introduce a distribution-guided Mapper algorithm named D-Mapper, that utilizes the property of the probability model and data intrinsic characteristics to generate density-guided covers and provides enhanced topological features. Moreover, we introduce a metric accounting for both the quality of overlap clustering and extended persistent homology to measure the performance of Mapper-type algorithms. Our numerical experiments indicate that the D-Mapper outperforms the classical Mapper algorithm in various scenarios. We also apply the D-Mapper to a SARS-COV-2 coronavirus RNA sequence dataset to explore the topological structure of different virus variants. The results indicate that the D-Mapper algorithm can reveal both the vertical and horizontal evolution processes of the viruses. Our code is available at https://github.com/ShufeiGe/D-Mapper.
Conclusion: The D-Mapper algorithm can generate covers from data based on a probability model. This work demonstrates the power of fusing probabilistic models with Mapper algorithms.
keywords:
Topology data analysis, Mapper, mixture model, Extended persistence
@@ 1 Introduction
Topology data analysis (TDA) is a specific area of investigating data shapes based on topological theory. Topology is a powerful tool to study the shape of objects by finding invariants that remain unchanged under continuous deformations. These invariants can reflect objects’ intrinsic features. TDA attempts to identify and utilize these topology invariants in data analysis. The Mapper algorithm is a simple but powerful TDA tool for visualizing and clustering data. The algorithm constructs a graph in which each vertex represents a cluster, and each edge represents the two adjacent clusters that share some elements. The colour and size of vertices are defined by users and can provide more information. For example, the size of vertices and the colour depth can represent the average value and the number of elements in the cluster. The Mapper algorithm was first proposed in [1] and has been applied in various domains, such as social media [2], fraud detection [3], and evolutionary computation [4]. The Mapper algorithm is especially suitable for biology data, which are usually complex and high-dimensional [5]. For example, the Mapper algorithm was applied to breast cancer transcriptional microarray data and successfully identified two subgroups with $100\%$ survival rate [6]. It is also used to analyze transcriptional programs that control cellular lineage commitment and differentiation during development, and the proposed scTDA identified four transient states over time [7]. Exploring the conformation space of proteins is another important task in computational biology, and the Mapper algorithm has been successfully applied to analyze intermediate conformations, the results are closely consistent with experimental findings [8].
@@
The classic Mapper algorithm requires a filter function $f:X\rightarrow\mathbb{R}^{p}$ to project data $X$ onto the Euclidean space $\mathbb{R}^{p}$ and a set of open cover $\mathcal{U}$ on the projection data.In most of the literature, $p$ is set to $1$, i.e. the filter function $f$ maps data from a high dimensional space to a real line. Different filter functions and open covers may result in different outputs, so it is necessary to select them carefully. Improper filter functions or open covers may not accurately reveal data shapes, resulting in poor overlap clustering. A filter function is chosen primarily based on features of interest and also depends on specific applications. Determining the optimal parameters of the model, such as the overlapping rates and interval lengths, usually requires extensive manual tuning and experiments. As pointed out in [9], cover choices lead to a broad range of possible Mapper outputs, resulting in many different graph depictions and clustering results. Unreasonable choices may cause the disappearance of certain topological features.
@@
Many works have been proposed to improve the performance of the classic Mapper. For example, [10] proposed to generate covers for the dataset by constructing a set of balls directly. This method saved the trouble of choosing the filter function. [11] used a fuzzy clustering algorithm to generate covers automatically with random overlapping ratios. To achieve adaptive cover construction, an information criteria was developed based on an X-means algorithm to generate adaptive covers [12].
@@
In this work, alternatively, we propose a distribution-guided Mapper algorithm to relax the restriction of regular covers and overlapping ratios. Our proposed algorithm utilizes the property of the probability model and data intrinsic characteristics to generate distribution-guided covers and provides enhanced topological features. Unlike the classic Mapper algorithm, our proposed D-Mapper does not rely on predefined static overlapping rates and interval lengths, but instead, we fit projected data to a mixture distribution model and generate flexible covers automatically. Moreover, the performance of the Mapper algorithm is hard to evaluate due to the complexity of the output graph structure. Methods in most of the literature only evaluate it from the clustering perspective and overlook the quality of topological structure [11, 12]. In this paper, we introduce a metric that can quantitatively and objectively evaluate the performance of the Mapper algorithm from both clustering and topological aspects. We use the extended persistent homology as a tool to capture topological signatures, and then we distinguish noise and signal on the diagram by constructing a confidence set on the diagram, the signal rate can be used as a metric to evaluate the quality of topological structure. The simulation studies indicate that the D-Mapper algorithm outperforms the classic Mapper algorithm. The remainder of the paper is organized as follows. We briefly review the classic Mapper algorithm and introduce our D-Mapper and evaluation metric in the section 2. Section 3 focuses on the comparison of the classic Mapper and the D-Mapper on simulated datasets and the application of the proposed D-Mapper algorithm on a real-world SARS-COV-2 RNA sequence dataset. We discuss and summarize our work in sections 4 and 5.
2 Methods
@@ 2.1 Basic notions
Before we introduce our method, we first briefly review the essential background of the Mapper algorithm. The theoretical foundation of the Mapper algorithm is the Nerve theorem which guarantees the nerves produced by a cover on a topological space $X$ is homotopy equivalent to that space. We introduce the following notions to describe the Nerve theorem [13, 14].
@@
Given a set $P=\{p_{0},...,p_{k}\}\subset\mathbb{R}^{p}$ of $k+1$ affinely independent points, a $k$-dimensional simplex $\sigma$, or $k$-simplex for short, spanned by $P$ is the set of convex combinations such that:
@@
b) The intersection of any two simplices of $K$ is either empty or a common face of the two.
The union of the simplices of $K$ constitutes the underlying space of $K$, denoted as $|K|$, which inherits from the topology space of $\mathbb{R}^{p}$. Thus the geometric simplicial complex can also be regarded as a topological space.
@@
Suppose $\mathcal{U}=(u_{i}),i\in I$ is a collection of open subset of a topological space $X$, then we say $\mathcal{U}$ is an open cover of $X$ if $X=\bigcup_{i\in I}u_{i}$.
Given an open cover of topological space $X$, $\mathcal{U}=(u_{i})$, the nerve of $\mathcal{U}$ is an abstract simplicial complex C$(\mathcal{U})$ with vertex set $\mathcal{U}$. With these definitions, we can introduce the most important theorem in constructing the Mapper algorithm.
@@
Let $\mathcal{U}=(u_{i}),i\in I$ be an open cover of a paracompact space $X$ by open sets such that the intersection of any sub-collection of the $u_{i}$’s is either empty or contractible. Then, $X$ and the nerve C($\mathcal{U}$) are homotopy equivalent.
This theorem allows us to map the topology of continuous into abstract combinatorial structures by building a nerve complex. It bridges the gap between continuous space and its discrete representation. This crucial theorem is also a motivation for the construction of the Mapper algorithm.
@@ 2.2 Mapper algorithm
The Mapper algorithm is an important tool to construct the discrete version of Reeb graphs which encode connected information of the support manifold [15]. Algorithm 1 depicts the classic Mapper algorithm. Firstly, the original data is projected onto a real line by a user-specified filter function (Algorithm 1, line 1). To construct a cover $\mathcal{U}$ on the projected data, the number of components $n$ and overlapping ratio $p$ should be chosen carefully (Algorithm 1, line 2). The improper choice of these parameters may lead to failure estimation of the shape of data. With these two parameters, a cover with $n$ regular intervals can be obtained (Algorithm 1, line 3). Then, the inverse image of the cover $\mathcal{U}$ is achieved on the original data, generating some hypercubes in the original data space. This process is often called pulling back (Algorithm 1, line 4). Finally, clustering data points within each hypercube. Each cluster corresponds to a vertex in the Mapper graph. If two vertices share any elements, an edge is added between these two vertices (Algorithm 1, line 5).
@@
1:Choose a proper filter function $f$ to project data onto a real line, $f:X\rightarrow\mathbb{R}$.
2:Choose a component number $n$ and overlapping percentage ratio $p$.
@@
4:Pull back the intervals of projected data to the original space, $f^{-1}(\mathcal{U})$.
@@
The classic Mapper algorithm is simple but powerful for exploring and visualizing data, while the selection of parameters $n,~{}p$ involves extensive manual tuning and the algorithm is sensitive to these parameters. In addition, the Mapper algorithm’s flexibility is often restricted by regular intervals and fixed overlapping ratios, which may hinder the discovery of complex data structures.
@@ 2.3 D-Mapper
The restriction of regular intervals with fixed overlapping ratios is one of the major limitations of the Mapper algorithm. We propose a distribution-guided Mapper algorithm to generate flexible covers to reflect the underlying data structures. Our proposed method automatically chooses the overlapping ratios based on the distribution of the projection data and produces more flexible covers to reveal the data shapes more accurately. The key idea of our algorithm is to fit the projected data with a mixture probability model. Each component in the mixture model can be viewed as an interval, and the probability (likelihood) of each data point assigned to each interval can be explicitly calculated. Once we get the mixture distribution of the projected data, we can create intervals based on the distribution in many ways.
Here we introduce a simply way to construct intervals based on the quantile $\alpha$ of the probability distribution. With proper selection, the $\alpha$ quantile interval of each component of the mixture model can automatically produce some overlaps. Figure 1 shows the idea of naturally producing intervals by quantile $\alpha$. This attribute provides a natural scheme for constructing flexible covers on the projected data. The specific procedures are as follows:
@@
Figure 1: An illustration of intervals produced by the D-Mapper algorithm. The deep blue line represents the probability density function of each component in the GMM. The shallow blue dashed line presents the probability density function of the GMM. The orange lines are intervals that produced naturally by symmetry quantile $\alpha$. (a) When $\alpha=0.01$, there are overlaps between adjacent intervals. (b) When $\alpha=0.1$, there is a gap between the first and second interval. $\alpha$ controls the overlap of intervals, it should be chosen carefully.
In both our simulation and real data experiments, we implement the Gaussian mixture models (GMM) to fit the projected data due to its simplicity and flexibility. The model inference is done via the expectation maximization (EM) algorithm [16]. Other mixture models or any distribution with multiple modes can be used as alternatives to the GMM.
By incorporating a mixture model into the Mapper algorithm, it can produce flexible intervals and overlaps. We call this distribution-guided Mapper algorithm as D-Mapper algorithm. With these flexible intervals, we can construct the output graph similar to the classic Mapper algorithm: pulling back these intervals to the original data space, then clustering data points within each interval separately, all sub-groups constitute the vertices of the nerve, and adding an edge between two vertices if there are any shared data points. Algorithm 2 gives details of the D-Mapper.
@@
1:Choose a proper filter function $f$ to project data onto a real line, $f:X\rightarrow\mathbb{R}$.
@@
3:Fit the projected data to a mixture model.
@@
10:Pull back the intervals of the projected data to the original space, $f^{-1}(\mathcal{U})$.
11:Cluster on the refined cover and build the nerve by clustering result.
The quantile $\alpha$ controls the overlapping ratios: larger $\alpha$ leads to lower overlapping ratios. Distributions of components in mixture models are often heterogeneous, resulting in flexible quantile intervals automatically. Proper $\alpha$ selection is crucial to ensure all points are covered. The Mapper algorithms usually require pairwise overlap (i.e. each interval has overlap with its neighbours)[15, 10, 11, 12]. For D-Mapper, the pairwise overlap property can also be preserved, as shown in Figure 1 (a). However, larger $\alpha$ may result in disjoint intervals, Figure 1 (b) gives an example of disjoint intervals caused by improper large $\alpha$. Thus, in this work, we propose a method to find the upper bound of $\alpha$ that guarantees the pairwise overlap property. By establishing this upper bound, we also significantly narrow the potential range of $\alpha$ values, thereby facilitating parameter tuning. Denote $F_{i}^{-1}$ the inverse of the cumulative density function of the $i$th ordered component, thus the $i$th interval $u_{i}$ is given by $[F_{i}^{-1}(\alpha/2),F_{i}^{-1}(1-\alpha/2)]$, $i=1,\ldots,n$. The idea of this method is to find the upper bound $\alpha^{\prime}$, that guarantees the intersection of each paired adjacent intervals is not empty, $u_{i}\cap u_{i+1}\neq\emptyset$, i.e., $F_{i}^{-1}(1-\alpha/2)\geq F_{i+1}^{-1}(\alpha/2)$, $i=1,2,\ldots,n-1$. Notice that although pairwise overlap is a good property for a Mapper graph, this property is not necessary for all situations. One can construct a good Mapper graph with some paired adjacent intervals being non-overlap as long as the union of intervals can cover all data points. As shown in Figure 6 (a) in Section 3, since there are no points between the two disjoint circles, disjoint paired adjacent intervals should be presented in the corresponding Mapper output graph. We add one threshold parameter $\alpha^{*}$ to allow for disjoint paired adjacent intervals (lines 6-10 of Algorithm 3). In practice, we suggest to set $\alpha^{*}=0.005$. Algorithm 3 describes how to find the upper bound of $\alpha$.
@@
6:     if $\alpha_{i}^{\prime}\geq\alpha^{*}$ then